From GlueXWiki
Jump to: navigation, search

September 15, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • First package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update
  4. Chamber testing at EEL126 (Beni)
  5. Other: latest pictures of the package


Participants: Eugene, Bill, Dave, Mark, Simon, Beni, Glenn, and Lubomir.


- Started stacking again the package; we have already installed the first cell and the first cathode of cell#2. The second cathode of cell#2 and the second wire plane are also ready for installation after replacing two wires (see explanations below). Cell#3: the wire plane is ready except that four broken resistors need to be replaced. The two cathodes for cell#3 were checked with a generator and a problem in one connector was found: attenuation of the signal on several channels; the conductive tape on that connector will be re-heated. On all the frames in this stack (except the bottom end window and wire frame) we put new O-rings, because we observed some gas leakage. All the cathodes for the first package are ready except one; it has been tensioned and the next operation is to cut the flaps.

- Beni asked if the problems with the conductive tape show up at the same place on the connector due to uneven pressure. No, but Bill proposed to install a manometer at the cylinder to monitor that the pressure that comes from the compressor outside. Regarding resistor breaking, we decided not to use lead bricks during the gluing of the PCB ring.

- Testing the first two cells in the package. The first cell that was fixed before it was tested again and all the channels were OK. Two problems on the second cell: one wire broken between the pad and the epoxy and a point source of noise inside the chamber was found. Not clear how did the wire broke since that place was covered with epoxy and then the wire was checked by the tension measurements. The noise showed up on one wire and on both cathodes, pointing to a fixed position; that's how it was identified. Both wires were replaced. It was difficult to conclude what was the source of the noise: Casey saw several kinks in that area. After the wire was removed we looked with a microscope and found a small fiber-like piece looped over the wire, but we were not sure if that was at the same place. To bubble to the two cells we had to increase the flow to 100 ccpm (compared to 20 for the first cell before) meaning some increased gas leakage.

- There was a power outage at Blue Crab on Monday that affected mostly the AC unit for several hours.

- 55Fe source: the actual activity of the source when it arrived was 28 micro-Ci according to the manufacturer. RadCon informed us about this yesterday and they said we should wait several months till we can use it at Blue Crab. RadCon will look if there's another source available for us. Eugene proposed to attenuate the source with some material, but after the meeting RadCon clarified that the limit is on the activity not on the intensity.


- Bill proposed a new design for the rods/pins used to assemble the package. The idea is to make their height adjustable depending on the number of cells that are in the stack. Everybody agreed that this will save effort and possible damage during assembly since the holes on the frames are very tight. The third gusset ring will be ready by the end of next week.


- Chris is in vacation; will work on the board that will bring the signals from the daughter cards outside for testing. While testing the resistors on the PCB's, Anatoly found that some percentage (~5-10%) of the resistors have lower values by 10-30%. These are current limiting resistor, but still Lubomir will discuss with Fernando if these need to be replaced.

Chamber testing at 126

- Beni started rearranging the space at 126. He swapped the electronics and the gas racks and added another rack with a second VXS crate. Now there's enough space around the racks.


- For the Lehman review next month Eugene needs new results from the production chambers. Beni explained his results about the chamber efficiency for the second cell. Lubomir will also prepare some plots from the test of the second cell using fAD125. As for the first production package, the plan is to have it ready before the review but most likely by that time still there will be no results from the tests with the DAQ.